OpenStudy (sanchez9457):

What's the difference between V_x = v_x + a_xt and Vavg_x = v_x + 1/2a_xt? you could only use them with constant acceleration but whats the difference?

OpenStudy (anonymous):

the first equation you have written is the solution of differential equation dv/dt=a . And the second equation you have written is the result that follows from the definition of average velocity for a uniformly accelerated motion. vf=vi+at.....................(1) where vf is the final speed vi is the initial speed average speed (for uniform as well as non uniform acceleration) is defined as v_avg = total distance travelled / total time taken for uniform acceleration total distance travelled = (vf^2 - vi^2) / 2a total time taken = (vf-vi) / a [from eqn 1] Hence v_avg = (vf+vi)/2 Eliminating vf by using eqn (1) we get v_avg = vi + at/2